Question.13

An electron enters a region of space containing a uniform 0.0000171-T magnetic field. Its speed is 109 m/s and it enters perpendicularly to the field. Under these conditions, the electron undergoes circular motion. Find the radius r of the electron\'s path, and the frequency f of the motion.

r:.....meter

f:....Hz
Answer #1

radius of electron r is given as

r = mv/qB   

where m = mass of electron = 9.1*10-31 kg , V = velocity = 109m/s , B =magnetic field = 0.0000171 T

q = charge of electron = 1.6*10-19

hence r = 36.25*10-15 m/sec

frequency f = qB/2*pi*m = 4.76*105 m/sec

r =